OpenStudy (anonymous):

What is the sum of the geometric sequence 1, –4, 16, … if there are 7 terms?

OpenStudy (callisto):

a=1 r=-4 n=7 use the formula in ur previous post and calculate the answer

OpenStudy (callisto):

r=-4/1 = -4

OpenStudy (callisto):

answer = 3277

OpenStudy (anonymous):

That's correct! = [( -4)^7 -1 ] / -5 = 3277